OpenStudy (anonymous):

prove the identity cos^2x+2/csc^2x=1+sin^2x

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ csc(\theta)=\cfrac{1}{sin(\theta)}}}\\ \quad \\ \quad \\ cos^2(x)+\cfrac{2}{csc^2(x)}=1+sin^2x\) replace the csc(x), and simplify